40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10929     Entry time: Thu Jan 22 03:21:24 2015
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Locks with large MICH offsets 

[Jenne, Diego, EricQ]

Tonight we worked on the acquisition sequence (including re-re-re-commissioning the UGF servos, hopefully for the last time...) for the PRFPMI with large MICH offsets. 

The procedure is all in the carm_up script, as far as things work.

We had some locklosses, but they were mostly due to non-carefulness on my part during the transitions between error signals, or the UGF servos getting upset because the oscillator peaks had gotten lost in the noise.  The one that I show here is our very last one of the night, where we are hitting the rails for the MICH signal, which is then causing the other loops to have to do weird things to try to compensate, and they lose lock.

Here also is a StripTool shot during that lock stretch.  I was in the middle of increasing the MICH offset to 75% of the fringe.  The yellow trace (called MICH_B_MON) is ASDC/POPDC normalized so that it always goes 0-1.  I was pleased to see that perhaps REFL11I and AS55Q are turning over, although as Q will tell us in a more detailed elog tomorrow, having a large MICH offset does weird things and moves the DARM zero-point.  So, maybe we aren't actually anywhere awesome yet. 

After some MICH offset, the maximum arm power is always going to be about 50, so arm powers of 8 or 10 equates to 100 pm.  We didn't get there tonight while on IR signals.


The locking sequence is now something like this:

  • Lock carm and darm on ALS, find resonances, move to 3 counts (roughly 3nm) offset.
  • Set PRMI up to acquire on REFL33I and ASDC/POPDC at 25% MICH fringe.  (After a while, I assume perhaps because the alignment is no longer tip-top, I have been by-hand reducing the MICH offset from -700counts which is 25% to -200counts, and then immediately putting it back to -700 after the PRMI acquires.)
  • Engage all 4 UGF servos
  • Reduce the CARM offset a bit, to 1.0 count, which gives arm powers of about 0.4 (with 50 being the max possible)
  • Transition CARM from ALS to sqrtInvTrans
  • Transition DARM from ALS to DC trans:  (TRY-TRX)/(TRX+TRY)
  • Reduce the oscillator amplitudes of the UGF servos
  • Reduce the CARM offset to powers of about 1
  • Ramp to 50% MICH fringe

After this, we tried a few times to lower the CARM offset, but kept losing lock, I think because the UGF servos went crazy.  The final lock, shown above, we lost because the MICH output was hitting the rails.

The problem with the MICH servo right now is the low SNR of the POPDC being used to normalize ASDC.  The control output is enormous, even if we have the 400Hz lowpass on.  We need to rethink our MICH servo, starting with a lower UGF, so that we're not injecting all this sensing noise all over the place. 


For tomorrow:

  • Re-look at MICH loop, to prevent sensing noise injection.
  • How does the large MICH offset affect our zero points for CARM and DARM?  Can we stay on DC transmission signals through 30 or 100 pm?
  • What to do next?  One or two of the locklosses were because the CARM detuned double cavity pole wasn't de-Q-ed enough, so still hit 0dB and created an unstable unity gain point.  Can we go to higher MICH offset, maybe 75%? 
  • Still need to figure out where our missing phase is for our LSC loops.  CARM and DARM are short on phase, and we could definitely use some more.  So, I will work on trying to give us filters that don't eat too much phase, but we still need to find that missing ~14 deg. 
ELOG V3.1.3-